# references.bib

### v2

@@ Abstract
For a graph family $\mathcal{F}$, let $\mathrm{ex}(n,\mathcal{F})$ and $\mathrm{spex}(n,\mathcal{F})$ denote the maximum number of edges and maximum spectral radius of an $n$-vertex $\mathcal{F}$-free graph, respectively, and let $\mathrm{EX}(n,\mathcal{F})$ and $\mathrm{SPEX}(n,\mathcal{F})$ denote the corresponding sets of extremal graphs. Wang, Kang, and Xue showed that if $\mathrm{EX}(n,\mathcal{F})$ consists of Turán graphs $T_{n,r}$ plus $O(1)$ edges, then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. Fang, Tait, and Zhai extended this result by showing if $e(T_{n,r})\leq\mathrm{ex}(n,\mathcal{F})<e(T_{n,r})+\lfloor n/2r\rfloor$ then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. In this paper we extend the result further and for many values of $r$ we can show that our result is best possible, answering a question of Fang, Tait, and Zhai.
@@ 1 Introduction
Thus, Theorem 1 shows that $c(r)\geq 1/2r$ for every $r$. In this paper we give improved lower and upper bounds on $c(r)$ when $r\geq 3$ and in many cases we can determine $c(r)$ exactly (see 1 below). When $r=2$, we prove that $c(2)\leq 1/3$ and conjecture that this is the correct value. These results also improve the upper bound $c(2)\leq 3/2$ obtained by considering the odd wheel $W_{15}$ [yuan2020extremal, cioabua2022spectral] (we are not aware of existing upper bounds for larger $r$).
@@ 1.1 Notation and definitions
We write $K_{n}$ for the complete graph on $n$ vertices; $K_{V_{1},\ldots,V_{r}}$ for the complete multipartite graph with partite sets $V_{1},\ldots,V_{r}$; $T_{n,r}$ for the $r$-partite Turán graph on $n$ vertices; $P_{\ell}$ for a path on $\ell$ vertices; and $M_{n}$ for a matching on $n$ vertices (or almost matching if $n$ is odd). For a graph family $\mathcal{F}$, let $\chi(\mathcal{F})=\min\{\chi(F):F\in\mathcal{F}\}$. Given a graph $G$, we write $V(G)$ and $E(G)$ for its vertex set and edge set, respectively; given a set $S\subseteq V(G)$, we write $N_{G}(S):=\{v\in V(G):v\sim u\text{ for all }u\in S\}$, or just $N(S)$ when the underlying graph is understood; $G[U]$ for the induced subgraph on $U$, $E[U]$ for $E(G[U])$ and $e(U)$ for $|E(U)|$; $G[U,W]$ for the bipartite subgraph induced by $U$ and $W$; $A(G)$ for the adjacency matrix of $G$ and $\lambda(G)$ for the spectral radius of $A(G)$. For a vertex $v\in V(G)$ and a set $S\subseteq V(G)$, we write $N_{S}(v):=N(v)\cap S$ and $d_{S}(v)$ for $|N_{S}(v)|$; we write $G-v$ for the graph obtained by deleting the vertex $v$; and $d^{\prime}(v)$ for the number of walks of length two starting from $v$. For $n\in\mathbb{N}$, let $U(G,n)$ be a graph on $n$ vertices consisting of as many disjoint copies of $G$ as possible with isolated vertices in the remainder. We write $G\cup H$ and $G+H$ for the disjoint union and join, respectively, of $G$ and $H$ and $k\cdot G$ for the union of $k$ disjoint copies of $G$.
@@ 1.2 Main results
Suppose that $r\geq 3$, $k\in\mathbb{N}$, and
$k>\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right).$
Then there is a finite family of graphs $\mathcal{F}$ such that $\chi(F)=r+1$, $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+\frac{k-1}{kr}n+O(1)$, and $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$ for an infinite sequence of $n$. Thus, $c(r)\leq\frac{k-1}{kr}.$
@@
If $r\geq 3$ and
$\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)$
is not an integer, then $c(r)=\frac{k-1}{kr}$, where $k$ is the ceiling of the expression above.
We illustrate 1 by giving some precise values of $c(r)$.
$r$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$
$c(r)$ $5/18$ $7/32$ $9/50$ $11/72$ $13/98$ $15/128$ $17/162$ $19/200$
Additionally, the asymptotics of $c(r)$ can be deduced from Theorem 2 and 1, namely $c(r)\sim 1/r$ as $r\to\infty$.
@@
We conjecture that a weaker result holds when $r=2$.
@@ 1.3 Applications
To obtain new applications of Theorem 2, we are interested in families $\mathcal{F}$ such that $e(T_{n,r})+\left\lfloor\frac{n}{2r}\right\rfloor\leq\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$. One example is the graph $P_{2p+1}^{p}$ obtained by joining all vertices at distance at most $p$ in a path of length $12p+$ 1, where $p\geq 3$. Then Theorem 1.2 of [yuan2022extremal] gives that $\mathrm{EX}(n,\mathcal{F})$ contains a graph obtained from $T_{n,p}$ with a matching embedded in one part, hence Theorem 2 implies that $\mathrm{SPEX}(n,P_{2p+1}^{p})\subseteq\mathrm{EX}(n,P_{2p+1}^{p})$ for $n$ large enough.
We are not aware of any other existing extremal results in this category. However, with the goal of constructing graphs $F$ for which we can show $\mathrm{ex}(n,F)=e(T_{n,r})+\frac{n}{2r}+O(1)$ we began exploring specific small graphs. Other than the path power discussed above, the simplest such $F$ we found is described as follows. The graph $F_{1}$ consists of two paths on four vertices $bcde$, $fghi$, with additional edges $bf,bg,cf,ch,dg,di,eh,ei$ and additionally one dominating vertex $a$.
@@
    If the added edges are $u_{1}v_{2}$ and $v_{1}v_{2}$ then we find a copy of $F_{1}$ by identifying $bf$ with $u_{1}u_{2}$ and $ei$ with $v_{1}v_{2}$.
@@
    If the added paths are $u_{1}u_{2}u_{3}$ and $v_{1}v_{2}$ then we find a copy of $F_{1}$ by identifying $gbc$ with $u_{1}u_{2}u_{3}$ and $ei$ with $v_{1}v_{2}$.
@@
    Suppose $F_{1}$ appears in this graph. Assume without loss of generality that the matching is embedded in $V_{1}$, $a\in V_{3}$, and $b\in V_{1}$. If $f\in V_{1}$ then $c,g\in V_{2}$ so $d,h\in V_{1}$ and so $e,i\in V_{2}$, contradicting $e\sim i$. If $f\in V_{2}$ then $c,g\in V_{1}$ so $gbc$ is a $P_{3}$ in $V_{1}$, a contradiction.
@@ 2.1 Structure of edge extremal graphs
  • $K_{1,m}+T_{m(r-1),r-1}$ contains a graph obtained by adding $m-1$ edges to $T_{m(r-1),r-1}$, so $Q(m-1)r\geq m$ by assumption and $Q\geq\left(1-\frac{1}{m}\right)/r$, which contradicts $Q<1/r$ for large enough $m$.
@@
The Erdős-Simonovits stability theorem gives the following lemma.
Lemma 3 (Erdős-Simonovits [erdos1967some],[erdos1966some],[simonovits1968method]).
Let $V(G)=V=V_{1}\cup\cdots\cup V_{r}$ where $|V_{i}|\in\{\lfloor n/r\rfloor,\lceil n/r\rceil\}$, and $G$ is obtained from $K_{V_{1},\ldots,V_{r}}$ by adding a graph $A$ and removing a graph $B$ with $e(A)+e(B)\leq\varepsilon n^{2}$. Let $X=\{v\in V:d_{B}(x)\geq 2\sqrt{\varepsilon}n\}$, and note that we have
@@
If we delete $v$ and add a $V$-ordinary vertex we obtain
@@
and if we delete $v$ and add an $V$-ordinary vertex we obtain
@@
By 4, we may define a new partition $V=W_{1}\sqcup\cdots\sqcup W_{r}$ as follows. If $v\in V_{i}-X$ then let $v\in W_{i}$. If $v\in X$ then let $v\in W_{i}$ where $i$ is chosen such that $d_{V_{i}}(v)\leq\varepsilon_{1}n$ and $d_{V_{j}}(v)\geq n/r-\varepsilon_{2}n$ for every $j\neq i$. Let $w_{i}:=|W_{i}|$. Note that $w_{i}\in(n/r-\sqrt{\varepsilon}n+O(1),n/r+\sqrt{\varepsilon}n+O(1))$. Thus for each $v\in W_{i}$, for each $j\neq i$, $d_{W_{j}}(v)\geq w_{j}-(\varepsilon_{2}+2\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{3}n$, since $\varepsilon_{2}+2\sqrt{\varepsilon}<\varepsilon_{3}$, (vi).
@@
Let $G^{\prime}$ be obtained from $G$ by deleting some vertex and adding an ordinary vertex $u\in W_{i}$. Then $G^{\prime}$ is $\mathcal{F}$-free.
@@
Let $v\in W_{i}^{o}$ not be ordinary. Since $d_{W_{i}}(v)=0$, replacing $v$ with an ordinary vertex increases the number of edges, which by 6 and $G\in\mathrm{EX}(n,\mathcal{F})$ is a contradiction. ∎
@@
If the tree components of $G[W_{1}]$ of order at most $k-1$ are $T_{1},\ldots,T_{j}$, then we have
@@
Every vertex in $W_{1}^{o}$ is $W$-ordinary.
@@
Below we refer to $H_{ij}$ for various graphs $H$ without restating the full setup when it is clear from context. We will make statements such as “$H$ satisfies $|w_{i}-w_{j}|\leq a$ for all $i,j$,” where the $w_{i}$ are assumed to be the values of $|W_{i}|$ which satisfy the assumptions of 13.
@@ 2.2 Structure of spectral extremal graphs
Let $m$ be a large enough even constant satisfying 16. Let $M=\max\{|V(F)|:F\in\mathcal{F}\}$. Suppose $G\in\mathrm{SPEX}(n,\mathcal{F})$, where $n$ is large enough. Choose positive constants $\varepsilon,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$,$\varepsilon_{4},\varepsilon_{5},\varepsilon_{6}$ satisfying the following inequalities (we do not attempt to avoid redundancy or optimize the constants).
@@
Suppose not. Note that an ordinary vertex satisfies $d(u)\geq(1-1/r-\sqrt{\varepsilon})n+O(1)$.
@@
By replacing $v$ with an ordinary vertex, we obtain that
@@
Then replacing $v$ by an ordinary vertex gives
@@
If the tree components of $G[W_{1}]$ of order at most $k-1$ are $T_{1},\ldots,T_{a}$, then we have
@@
$00\geq 3\left(1-\frac{1}{r}\right)-2+\frac{1}{r}+o(1)\geq 1-\frac{2}{r}+o(1)>$ 0
@@
$00\geq 3\left(1-\frac{1}{r}\right)-1-\frac{2}{r}+o(1)\geq 2-\frac{5}{3}+o(1)>$ 0
@@
Choose a non-pendant vertex $u$ as the root of $T$ and choose $v\in V(T)$ at maximum distance from $u$. Since $T$ is not a star, $d(v,u)\geq 2$ and also $v$ is a pendant vertex. Let $v\sim_{T}w$. Since $v$ is at maximum distance from $u$ and $w\neq u$, $w$ has a unique neighbor $w^{\prime}$ such that $w^{\prime}$ is not a pendant vertex. Let $T^{\prime}$ be obtained by deleting the edge $wv^{\prime}$ and adding the edge $w^{\prime}v^{\prime}$ for every pendant $v^{\prime}\sim_{T}w^{\prime}$, and note that $T^{\prime}$ has more pendant vertices than $T$. Now $K^{\prime}$ is obtained from $K$ by performing a Kelmans transformation and so $\lambda(K^{\prime})\geq\lambda(K)$ (see e.g. [csikvari2009conjecture]). ∎
@@
Observe that the number of walks of length two in a path of order $k\geq 4$ is $64k-$ 6 and the number of walks of length two in $K_{1,k-1}$ is $k(k-1)$. Therefore,
@@
Now, the assumption $Q<c_{1}(r)$ and $1\leq k\leq 1/(1-Qr)$ imply that $02-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}>$ 0. Since $k>4/3$, we have $\frac{4(k-1)}{kr}-\frac{2}{r}>-\frac{4(k-1)}{kr}$ and so the bracketed expression above is positive, giving a contradiction. ∎
@@
As stated above, the assumption $Q<c_{1}(r)$ gives $02-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}>$ 0, and we will use this estimate below. Let $H\in\mathrm{EX}(n,\mathcal{F})$, with partition $V_{1}\sqcup\cdots\sqcup V_{r}$ according to 1, and $v_{i}:=|V_{i}|$. If $v_{1}<v_{i}$ for some $i$ then $e(H_{i1})=e(H)$, so we may assume without loss of generality that $v_{1}\geq v_{i}$ for every $i$. Then either $w_{1}=v_{1}$ or $w_{1}=v_{1}-1$. Assume for a contradiction that $e(H)>e(G)$.
@@
Since $02-\frac{k-5+6/k}{r-1}>$ 0, this gives a contradiction.
Case 2: $w_{1}=v_{1}-1.$ Then without loss of generality, $w_{2}=w_{1}+1=v_{1}$. Let $G^{\prime}$ be obtained from $G$ by replacing every vertex in $W^{m}$ with an ordinary vertex. Let $a=e(G^{\prime})-e(G)=O(1)$.
@@
Let $u^{\prime}$ be the ordinary vertex in $W_{2}$ which was transferred to obtain $G_{21}^{\prime}$. Let $\mu=\lambda(G^{\prime}),\rho=\lambda(G_{21}^{\prime})$, with corresponding eigenvectors $y,z$, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in W_{1}\cup\{u^{\prime}\}}z_{v}=1$. Let $w_{2},\ldots,w_{h}=w_{1}+1$ and $w_{h+1},\ldots,w_{r}=w_{1}$. Similarly to 44, we obtain the following preliminary estimates. First, $\rho-\mu=O(n^{-1})$. For $v\in W_{1}$,

### v3

# A sharp spectral extremal result for general non-bipartite graphs

@@ Abstract
For a graph family $\mathcal{F}$, let $\mathrm{ex}(n,\mathcal{F})$ and $\mathrm{spex}(n,\mathcal{F})$ denote the maximum number of edges and maximum spectral radius of an $n$-vertex $\mathcal{F}$-free graph, respectively, and let $\mathrm{EX}(n,\mathcal{F})$ and $\mathrm{SPEX}(n,\mathcal{F})$ denote the corresponding sets of extremal graphs. Wang, Kang, and Xue showed that if $r\geq 2$ and $\mathrm{ex}(n,F)=e(T_{n,r})+O(1)$ then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. Fang, Tait, and Zhai extended this result by showing if $e(T_{n,r})\leq\mathrm{ex}(n,\mathcal{F})<e(T_{n,r})+\lfloor n/2r\rfloor$ then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough, and asked for the maximum constant $c(r)$ such that $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+(c(r)-\varepsilon)n$ guarantees such containment. In this paper we determine $c(r)$ exactly for all $r\geq 3$.
Keywords: Spectral Turán problem, Turán graph
2020 MSC: 05C35, 05C50
@@ 1 Introduction
Thus, Theorem 1 shows that $c(r)\geq 1/2r$ for every $r$. In this paper we determine $c(r)$ exactly for every $r\geq 3$. When $r=2$, we prove that $c(2)\leq 1/3$ and conjecture that this is the correct value. These results also improve the upper bound $c(2)\leq 3/2$ obtained by considering the odd wheel $W_{15}$ [yuan2020extremal, cioabua2022spectral] (we are not aware of existing upper bounds for larger $r$).
@@ 1.1 Notation and definitions
We write $K_{n}$ for the complete graph on $n$ vertices; $K_{V_{1},\ldots,V_{r}}$ for the complete multipartite graph with partite sets $V_{1},\ldots,V_{r}$; $T_{n,r}$ for the $r$-partite Turán graph on $n$ vertices; $P_{\ell}$ for a path on $\ell$ vertices; and $M_{n}$ for a matching on $n$ vertices (or almost matching if $n$ is odd). For a graph family $\mathcal{F}$, let $\chi(\mathcal{F})=\min\{\chi(F):F\in\mathcal{F}\}$. Given a graph $G$, we write $V(G)$ and $E(G)$ for its vertex set and edge set, respectively; given a set $S\subseteq V(G)$, we write $N_{G}(S):=\{v\in V(G):v\sim u\text{ for all }u\in S\}$, or just $N(S)$ when the underlying graph is understood; $G[U]$ for the induced subgraph on $U$, $E[U]$ for $E(G[U])$ and $e(U)$ for $|E(U)|$; $G[U,W]$ for the bipartite subgraph induced by $U$ and $W$; $E(U,W)$ for the set of edges with one endpoint in $U$ and one endpoint in $W$ (where possibly $U\cap W\neq\emptyset$); $A(G)$ for the adjacency matrix of $G$ and $\lambda(G)$ for the spectral radius of $A(G)$. For a vertex $v\in V(G)$ and a set $S\subseteq V(G)$, we write $N_{S}(v):=N(v)\cap S$ and $d_{S}(v)$ for $|N_{S}(v)|$; we write GvG-v for the graph obtained by deleting the vertex $v$; and $d^{\prime}(v)$ for the number of walks of length two starting from $v$. For $n\in\mathbb{N}$, let $U(G,n)$ be a graph on $n$ vertices consisting of as many disjoint copies of $G$ as possible with isolated vertices in the remainder. We write $G\cup H$ and $G+H$ for the disjoint union and join, respectively, of $G$ and $H$ and $k\cdot G$ for the union of $k$ disjoint copies of $G$.
@@ 1.2 Main results
Suppose that $r\geq 3$, $k\in\mathbb{N}$, and $k>\alpha$. Then there is a finite family of graphs $\mathcal{F}$ such that $\chi(F)=r+1$, $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+\frac{k-1}{kr}n+O(1)$, and $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$ for an infinite sequence of $n$. Thus, $c(r)\leq\frac{k-1}{kr}.$
Combining Theorem 2 and 1 allows us to determine $c(r)$ for $r\geq 3$:
@@
For every $r\geq 3$, we have
$c(r)=\frac{\lceil\alpha\rceil-1}{\lceil\alpha\rceil r}.$
As a consequence of 1, we have $c(r)\sim 1/r$ as $r\to\infty$.
@@
Given Theorem 1 and 2 one can show that $c(2)\in\{1/4,1/3\}$. We conjecture that $c(2)=1/3$:
@@ 1.3 Applications
To obtain new applications of Theorem 2, we are interested in families $\mathcal{F}$ such that $e(T_{n,r})+\left\lfloor\frac{n}{2r}\right\rfloor\leq\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$. One example is the graph $P_{2p+1}^{p}$ obtained by joining all vertices at distance at most $p$ in a path of length $2p+1$, where $p\geq 3$. Then Theorem 1.2 of [yuan2022extremal] gives that $\mathrm{EX}(n,\mathcal{F})$ contains a graph obtained from $T_{n,p}$ with a matching embedded in one part, hence Theorem 2 implies that $\mathrm{SPEX}(n,P_{2p+1}^{p})\subseteq\mathrm{EX}(n,P_{2p+1}^{p})$ for $n$ large enough.
We are not aware of any other existing extremal results in this category. However, with the goal of constructing graphs $F$ for which we can show $\mathrm{ex}(n,F)=e(T_{n,r})+\frac{n}{2r}+O(1)$ we began exploring specific small graphs. Other than the path power discussed above, the simplest such $F$ we found is described as follows. The graph $F_{1}$ consists of two paths on four vertices bcdebcde, fghifghi, with additional edges $bf,bg,cf,ch,dg,di,eh,ei$ and additionally one dominating vertex $a$.
@@
    If the added edges are $u_{1}v_{2}$ and $v_{1}v_{2}$ then we find a copy of $F_{1}$ by identifying bfbf with $u_{1}u_{2}$ and eiei with $v_{1}v_{2}$.
@@
    If the added paths are $u_{1}u_{2}u_{3}$ and $v_{1}v_{2}$ then we find a copy of $F_{1}$ by identifying gbcgbc with $u_{1}u_{2}u_{3}$ and eiei with $v_{1}v_{2}$.
@@
    Suppose $F_{1}$ appears in this graph. Assume without loss of generality that the matching is embedded in $V_{1}$, $a\in V_{3}$, and $b\in V_{1}$. If $f\in V_{1}$ then $c,g\in V_{2}$ so $d,h\in V_{1}$ and so $e,i\in V_{2}$, contradicting $e\sim i$. If $f\in V_{2}$ then $c,g\in V_{1}$ so gbcgbc is a $P_{3}$ in $V_{1}$, a contradiction.
@@ 2.1 Structure of edge extremal graphs
  • $K_{1,m}+T_{m(r-1),r-1}$ contains a graph obtained by adding m1m-1 edges to $T_{m(r-1),r-1}$, so $Q(m-1)r\geq m$ by assumption and $Q\geq\left(1-\frac{1}{m}\right)/r$, which contradicts $Q<1/r$ for large enough $m$.
@@
We will rely on the following stability theorem of Erdős and Simonovits.
Lemma 3 (Erdős [erdos1967some, erdos1966some], Simonovits [simonovits1968method]).
Suppose that $\mathcal{F}$ is a finite family of graphs with $\chi(\mathcal{F})=r+1.$ For all $\varepsilon>0$ there exists $\delta>0$ such that the following holds: if $n$ is large enough, then any $n$-vertex $\mathcal{F}$-free graph $G$ with more than $e(T_{n,r})-\delta n^{2}$ edges can be obtained from $T_{n,r}$ by deleting and adding at most $\varepsilon n^{2}$ edges.
With respect to our graph $G\in\mathrm{EX}(n,\mathcal{F})$, and the constant $\varepsilon$ chosen above, this implies that $G$ can be obtained from $T_{n,r}$ by adding and deleting at most $\varepsilon n^{2}$ edges. Let $V(G)=V=V_{1}\sqcup\cdots\sqcup V_{r}$ where $|V_{i}|\in\{\lfloor n/r\rfloor,\lceil n/r\rceil\}$, and $G$ is obtained from $K_{V_{1},\ldots,V_{r}}$ by adding a graph $A$ and removing a graph $B$ with $e(A)+e(B)\leq\varepsilon n^{2}$. Let $X=\{v\in V:d_{B}(x)\geq 2\sqrt{\varepsilon}n\}$, and note that we have
@@
If we delete $v$ and add a $V$-ordinary vertex then we obtain a graph $G^{\prime}$ such that
@@
and if we delete $v$ and add an $V$-ordinary vertex then we obtain a graph $G^{\prime}$ such that
@@
By 4, we may define a new partition $V=W_{1}\sqcup\cdots\sqcup W_{r}$ as follows. If $v\in V_{i}-X$ then let $v\in W_{i}$. If $v\in X$ then let $v\in W_{i}$ where $i$ is chosen such that $d_{V_{i}}(v)\leq\varepsilon_{1}n$ and $d_{V_{j}}(v)\geq n/r-\varepsilon_{2}n$ for every $j\neq i$. Let $w_{i}:=|W_{i}|$. Note that $w_{i}\in(n/r-\sqrt{\varepsilon}n+O(1),n/r+\sqrt{\varepsilon}n+O(1))$. Thus for each $v\in W_{i}$, for each $j\neq i$, $d_{W_{j}}(v)\geq w_{j}-(\varepsilon_{2}+2\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{3}n$, since $\varepsilon_{2}+2\sqrt{\varepsilon}<\varepsilon_{3}$, (vi). Recall that $m$ is a large even integer satisfying 2.
@@
Let $G^{\prime}$ be obtained from $G$ by deleting some vertex and adding a $W$-ordinary vertex $u\in W_{i}$. Then $G^{\prime}$ is $\mathcal{F}$-free.
@@
Let $v\in W_{i}^{o}$ not be $W$-ordinary. Since $d_{W_{i}}(v)=0$, replacing $v$ with a $W$-ordinary vertex increases the number of edges, which by 6 and $G\in\mathrm{EX}(n,\mathcal{F})$ is a contradiction. ∎
@@
If the tree components of $G[W_{1}]$ of order at most k1k-1 are $T_{1},\ldots,T_{j}$, then we have
@@
Every vertex in $W_{1}^{o}$ is adjacent to all vertices in $V-W_{1}$.
@@
Below we refer to $H_{ij}$ for various graphs $H$ without restating the full setup when it is clear from context. We will make statements such as “$H$ satisfies $|w_{i}-w_{j}|\leq a$ for all i,ji,j,” where the $w_{i}$ are assumed to be the values of $|W_{i}|$ which satisfy the assumptions of 13.
@@ 2.2 Structure of spectral extremal graphs
Let $m$ be a large enough even constant satisfying 16. Let $M=\max\{|V(F)|:F\in\mathcal{F}\}$. Suppose $G\in\mathrm{SPEX}(n,\mathcal{F})$, where $n$ is large enough. Choose positive constants $\varepsilon,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$$\varepsilon_{4},\varepsilon_{5},\varepsilon_{6}$ satisfying the following inequalities (we do not attempt to avoid redundancy or optimize the constants).
@@
Suppose not. Note that a $V$-ordinary vertex satisfies $d(u)\geq(1-1/r-\sqrt{\varepsilon})n+O(1)$.
@@
By replacing $v$ with a $V$-ordinary vertex, we obtain that
@@
Then replacing $v$ by a $V$-ordinary vertex gives
@@
If the tree components of $G[W_{1}]$ of order at most k1k-1 are $T_{1},\ldots,T_{a}$, then we have
@@
$0\geq 3\left(1-\frac{1}{r}\right)-2+\frac{1}{r}+o(1)\geq 1-\frac{2}{r}+o(1)>0$
@@
$0\geq 3\left(1-\frac{1}{r}\right)-1-\frac{2}{r}+o(1)\geq 2-\frac{5}{3}+o(1)>0$
@@
Choose a non-pendant vertex $u$ as the root of $T$ and choose $v\in V(T)$ at maximum distance from $u$. Since $T$ is not a star, $d(v,u)\geq 2$ and also $v$ is a pendant vertex. Let $v\sim_{T}w$. Since $v$ is at maximum distance from $u$ and $\neq u, ww$ has a unique neighbor $w^{\prime}$ such that $w^{\prime}$ is not a pendant vertex. Let $T^{\prime}$ be obtained by deleting the edge $wv^{\prime}$ and adding the edge $w^{\prime}v^{\prime}$ for every pendant $v^{\prime}\sim_{T}w^{\prime}$, and note that $T^{\prime}$ has more pendant vertices than $T$. Now $K^{\prime}$ is obtained from $K$ by performing a Kelmans transformation and so $\lambda(K^{\prime})\geq\lambda(K)$ (see e.g. [csikvari2009conjecture]). ∎
@@
Observe that the number of walks of length two in a path of order $k\geq 4$ is 4k64k-6 and the number of walks of length two in $K_{1,k-1}$ is k(k1)k(k-1). Therefore,
@@
Now, the assumption $Q<c_{1}(r)$ and $1\leq k\leq 1/(1-Qr)$ imply that $2-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}>0$. Since $k>4/3$, we have $\frac{4(k-1)}{kr}-\frac{2}{r}>-\frac{4(k-1)}{kr}$ and so the bracketed expression above is positive, giving a contradiction. ∎
@@
As stated above, the assumption $Q<c_{1}(r)$ gives $2-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}>0$, and we will use this estimate below. Let $H\in\mathrm{EX}(n,\mathcal{F})$, with partition $V_{1}\sqcup\cdots\sqcup V_{r}$ according to 1, and $v_{i}:=|V_{i}|$. If $v_{1}<v_{i}$ for some $i$ then $e(H_{i1})=e(H)$, so we may assume without loss of generality that $v_{1}\geq v_{i}$ for every $i$. Then either $w_{1}=v_{1}$ or $w_{1}=v_{1}-1$. Assume for a contradiction that $e(H)>e(G)$.
@@
Since $2-\frac{k-5+6/k}{r-1}>0$, this gives a contradiction.
Case 2: $w_{1}=v_{1}-1.$ Then without loss of generality, $w_{2}=w_{1}+1=v_{1}$. Let $G^{\prime}$ be obtained from $G$ by replacing every vertex in $W^{m}$ with a $W$-ordinary vertex. Let $a=e(G^{\prime})-e(G)=O(1)$.
@@
Let $u^{\prime}$ be the $W$-ordinary vertex in $W_{2}$ which was transferred to obtain $G_{21}^{\prime}$. Let $\mu=\lambda(G^{\prime}),\rho=\lambda(G_{21}^{\prime})$, with corresponding eigenvectors $y,z$, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in W_{1}\cup\{u^{\prime}\}}z_{v}=1$. Let $w_{2},\ldots,w_{h}=w_{1}+1$ and $w_{h+1},\ldots,w_{r}=w_{1}$. Similarly to 44, we obtain the following preliminary estimates. First, $\rho-\mu=O(n^{-1})$. For $v\in W_{1}$,